OpenStudy (loser66):

Does n/2 divides n! Please, explain

OpenStudy (rational):

is n even ?

OpenStudy (loser66):

yes

OpenStudy (rational):

then yes because 1,2,3,...n/2,....n all of them divide n!

OpenStudy (rational):

n! = 1.2.3....n/2...n

OpenStudy (loser66):

Thank you!!

OpenStudy (loser66):

let me rearrange my question :)

OpenStudy (rational):

okay :) it is easy to see that `n!` is divisible by all numbers `<= n`

OpenStudy (rational):

*all positive integers

OpenStudy (loser66):

if n =9, then 4*5 divides 9!?

OpenStudy (loser66):

hahaha... my stupid question, of course it is ...

OpenStudy (loser66):

but how to arrange it in term of n??

OpenStudy (loser66):

I mean 4 = (n-1)/2 5 = (n+1) /2

OpenStudy (rational):

even numbers : 2k odd numbers : 2k-1

OpenStudy (rational):

if n is odd, (n-1) would be even, yes ?

OpenStudy (loser66):

yes

OpenStudy (rational):

you want to ask : Does (n-1)/2 divide n! when n is odd ?

OpenStudy (loser66):

Ok, let me retype it: Let S_n is a sequence If n is even, then Sn can be broken into 2-equal- parts, each of them has n/2 term I have to prove that l.c.m ( n/2, n/2) divides n!. The case close by your help. If n is odd, then Sn can be broken into 2 parts also, the first part has (n-1)/2 terms, the second part has (n+1)/2 terms. I have to prove \(l.c.m~~ (\dfrac{n-1}{2}, \dfrac{n+1}{2})\) divides n!

OpenStudy (loser66):

Surely, one is even and one is odd, so that \(l.c.m= (\dfrac{n-1}{2})( \dfrac{n+1}{2})\)

OpenStudy (loser66):

@rational I am sorry, the net is so bad here....

OpenStudy (rational):

when n is odd : n! = 1.2.3...(n-1)/2.(n+1)/2....n

OpenStudy (rational):

(n-1)/2 and (n+1)/2 are two DIFFERENT numbers less than n, so their product divides n! for obvious reasons ?

OpenStudy (loser66):

Yes, I got it. Thank you so much.